OpenStudy (anonymous):

What lines are paralle to the line whose equation is y=3x+5

OpenStudy (dobby1):

is it asking for the equation to a line that is parallel to the line of y=3x+5

OpenStudy (anonymous):

yes

OpenStudy (dobby1):

ok so here are some equations you can use y=3x+6, y=3x+4, y=3x+3,y=3x+2 etc

OpenStudy (dobby1):

@zombie1234

OpenStudy (anonymous):

thank you :)